### v1

# Dynamic-Dark SLAM: RGB-Thermal Cooperative Robot Vision Strategy for Multi-Person Tracking in Both Well-Lit and Low-Light Scenes

Abstract
In robot vision, thermal cameras have significant potential for recognizing humans even in complete darkness. However, their application to multi-person tracking (MPT) has lagged due to data scarcity and difficulties in individual identification. In this study, we propose a cooperative MPT system that utilizes co-located RGB and thermal cameras, using pseudo-annotations (bounding boxes + person IDs) to train RGB and T trackers. Evaluation experiments demonstrate that the T tracker achieves remarkable performance in both bright and dark scenes. Furthermore, results suggest that a tracker-switching approach using a binary brightness classifier is more suitable than a tracker-fusion approach for information integration. This study marks a crucial first step toward “Dynamic-Dark SLAM,” enabling effective recognition, understanding, and reconstruction of individuals, occluding objects, and traversable areas in dynamic environments, both bright and dark.
1 Introduction
Thermal cameras (T-cameras) hold great potential for person detection and tracking in dark or low-light scenes, representing a promising new application in robot vision. Despite the increasing demand for advanced vision systems that consider privacy, such as those used in nighttime festivals and security patrols, T-vision remains underutilized, particularly in its application to multi-person tracking (MPT). MPT is a crucial technology in robot vision and is employed in various applications, including user tracking [1], patrolling [2], dynamic SLAM [3], structure-from-occlusion [4], and traversability analysis [5]. Research on MPT has primarily been based on RGB vision [6], while challenges specific to T-vision remain largely unaddressed.
One of the major obstacles in applying T-vision to MPT is the lack of annotated datasets (bounding boxes + person IDs), which makes supervised training of a T-tracker infeasible. These datasets are typically represented as time-series pairs of bounding boxes and person IDs. However, most existing datasets focus on RGB vision, and thermal image datasets are not yet well established. Furthermore, person ID identification in thermal images is difficult and requires high-cost annotation efforts to achieve accuracy. As a result, applying tracking techniques designed for RGB vision to T-vision leads to a decrease in performance (Fig. 2). Notably, the patterns in thermal images heavily depend on factors such as clothing, making them more complex than traditional RGB image patterns. Due to this dependency, generalizing thermal image datasets to arbitrary environments is difficult, necessitating the development of T-trackers that do not rely on annotated datasets.
Figure 1: Crowded scenes and RGB-T cooperative vision.
In this study, we propose a novel cooperative MPT system that utilizes an RGB camera and a thermal (T) camera placed in the same pose, enabling effective tracking in both bright and dark conditions. The key idea is to train a T-tracker using knowledge transfer from an RGB-tracker in bright conditions. Among all possible combinations of (1) two scene types (Bright/Dark), (2) two modality types (RGB/T), (3) three tracker types (RGB/T/RGBT tracker), we exclude those that are clearly ineffective or redundant, leaving four viable combinations, as discussed in Section 4, which we then evaluate experimentally. The results demonstrate that the T-tracker achieves remarkable performance in both bright and dark scenes. Furthermore, our findings suggest that a tracker-switching approach using a binary brightness classifier is more suitable than a tracker-fusion approach for integrating the two trackers.
This study builds on previous research on teacher-student knowledge transfer [7], multi-person tracking under occlusion [8], and the recognition, understanding, and reconstruction of people, occlusions, and traversable areas [9], effectively integrating thermal SLAM [10] and dynamic SLAM [3]. This marks a crucial first step toward “Dynamic-Dark SLAM,” enabling the effective recognition, understanding, and reconstruction of individuals, occluding objects, and traversable areas in both bright and dark dynamic environments.
Figure 2: Output of the RGB tracker for the RGB image (left) and the T image (right).
2 Related Work
This study is related to several areas of research in robotic vision. In the field of robotics, research on multi-person tracking (MPT) has progressed using laser-based [11] and vision-based [12] methods, with the study of SLAM considering interactions between dynamic and static objects [3] being the most relevant to our work. However, most existing vision-based methods focus on RGB vision, leaving challenges specific to thermal (T) vision largely unresolved.
In the field of computer vision, studies on multi-sensor information fusion [13], such as RGB-LiDAR fusion, as well as research on teacher-student knowledge transfer [14], have been conducted. However, many of these studies rely on predefined cooperative schemes [15] where the teacher and student sensors are used complementarily at the same time, with little exploration of diverse cooperative strategies.
RGB-T fusion tracking has been studied in both non-deep approaches [16, 17, 18, 19] and deep approaches, including adaptive fusion [20], sparse fusion representation [21], cross-modal ranking [22], and graph models [23], as well as benchmark studies [24]. The approach of applying an RGB tracker [25] to the RGB-T tracking problem was explored in [26]. Our approach is based on the state-of-the-art [ByteTrack] tracker and differs in that it handles Dark scenes, where RGB trackers completely fail.
3 RGB-T Cooperative MPT
3.1 Formulation of MPT
MPT is formulated as a problem where a tracker $f$ takes an image sequence $x_{1:T}$ as input and outputs a sequence of tracking results $y_{1:T}$. Here, $x_{1:T}=(x_{1},\cdots,x_{T})$ represents the sequence of camera images, where $x_{t}$ denotes the image at time $t$. The tracking result is represented as $y_{t}=\{y_{t}[i]\}_{i=1}^{I_{t}}$, where $I_{t}$ is the number of bounding boxes detected in the image $x_{t}$ and each $y_{t}[i]$ consists of $(y_{t}^{\text{ID}}[i],y_{t}^{\text{BB}}[i],y_{t}^{\text{SCORE}}[i])$. Specifically, $y_{t}^{\text{ID}}[i]$ is the predicted ID of the person in the $i$-th bounding box, $y_{t}^{\text{BB}}[i]$ represents the predicted location and size of the $i$-th bounding box, and $y_{t}^{\text{SCORE}}[i]$ indicates the confidence score.
Given an annotated image sequence $(x_{1:T},y_{1:T})$, the tracker is trained as follows:
$\theta=\arg\min_{\theta}\mathcal{L}(f_{\theta}(x_{1:T}),y_{1:T})$ (1)
where $\mathcal{L}$ represents the loss function that represents the error between the predicted results and the annotations.
3.2 ByteTrack
In this study, we propose using ByteTrack[27] as the baseline tracker, which separates detection from tracking and remains robust even under temporary full or partial occlusion. Many existing trackers discard detection results with scores below a threshold, whereas ByteTrack adopts a simple and general approach that utilizes low-score objects. Its core module consists of an object detector that outputs bounding boxes, scores, and classes for each frame, as well as a Byte module that associates tracked object IDs with detection results. ByteTrack employs a Kalman filter to predict the position of objects in frame $t$ based on tracking information up to frame $t-1$. Then, it associates predicted bounding boxes with actual detections in frame $t$ based on Intersection over Union (IoU), matching the pairs with the highest IoU and selecting them based on a threshold.
This simple yet effective two-stage IoU matching method enables stable and fast tracking. In practice, ByteTrack demonstrated superior performance in our study. In addition to the main experiments, we conducted a comparative evaluation of three state-of-the-art trackers: OC-SORT [28], StrongSORT [29], and ByteTrack [27]. The RGB image sequence captured in a well-lit scene was used for this comparison because of the availability of high-quality ground-truth for such sequences. The results showed that ByteTrack significantly outperformed OC-SORT and slightly more stable than StrongSORT.
The baseline tracker ByteTrack was trained as follows:
$\theta^{\text{RGB}}=\arg\min_{\theta}\mathcal{L}^{\text{ByteTrack}}(f_{\theta}(x^{\text{RGB}}_{1:T}),y_{1:T})$ (2)
where $\mathcal{L}^{\text{ByteTrack}}$ represents the loss function used in ByteTrack. The training set $(x^{\text{RGB}}_{1:T},y_{1:T})$ consists of five RGB image datasets: MOT17[30], MOT20[31], CrowdHuman[32], CityPersons[33], and ETHZ[34]. These datasets encompass various environments and scenarios, including urban settings, high-density crowds, occlusion scenarios, diverse weather conditions, and long-range pedestrian tracking, enhancing the model’s adaptability.
3.3 Cooperative Strategy
An RGB camera and a thermal (T) camera were placed at the same position and synchronized. By positioning the lenses of both cameras as closely as possible, we minimized the misalignment of the fields of view, reducing coordinate errors between the RGB and thermal images. Additionally, temporal discrepancies and resolution differences were corrected after capturing the data to precisely align the coordinates between the RGB and thermal images.
The input image at time $t$ is defined as $x_{t}=(x_{t}^{\text{RGB}},x_{t}^{\text{T}})$, where $x_{t}^{\text{RGB}}$ represents the RGB image, and $x_{t}^{\text{T}}$ represents the thermal image at time $t$.
The optimal coordinate transformation between the two camera coordinate systems may vary slightly depending on the environment. For instance, in a crowded environment such as that shown in Figure 2, pedestrians move while avoiding obstacles, leading to motion trajectories constrained by the obstacles. Therefore, it is possible to optimize the coordinate transformation based on pedestrian movement trajectories. In this study, a small set of image pairs of the same pedestrians observed from different cameras was used as training data to adjust the coordinate transformation parameters, minimizing transformation errors. Since the experimental environment ensured that both cameras captured the same plane or had minimal parallax, we modeled the relationship between the cameras using a homography transformation.
The experiments were conducted in an indoor environment with obstacles such as desks and chairs. Furthermore, we evaluated the multi-person tracking performance in a scene with four pedestrians.
In the experiment, an iPhone 15 was used as the RGB camera, while a FLIR One Pro was used as the thermal camera. The iPhone 15 captured images at a resolution of 1920$\times$1080 at 30 fps, while the FLIR One Pro captured images at a resolution of 160$\times$120 at 8.7 fps. To clearly visualize the temperature distribution, we used the “grayscale mode” to enhance the visibility of the target objects.
Figure 3: Improved tracking results from the proposed method.
Using this synchronized dataset, tracking results (bounding boxes, person IDs, confidence scores) were obtained from the teacher model and used as pseudo-labels for thermal images.
$\bar{y}_{t}=f_{\theta^{\text{RGB}}}(x_{t}^{\text{RGB}})$ (3)
The student model adapted to thermal images by minimizing a loss function considering both positional and classification errors between its predictions and the teacher model’s tracking results.
$\theta^{\text{T}}=\arg\min_{\theta}\mathcal{L}^{\text{ByteTrack}}(f_{\theta}(x^{\text{T}}_{1:T}),\bar{y}_{1:T})$ (4)
Furthermore, a fusion model was introduced to complement the characteristics of both the teacher and student models.
$\hat{y}_{t}=f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{t}^{\text{RGB}},x_{t}^{\text{T}})$ (5)
where $f_{\theta^{\text{RGB}}\theta^{\text{T}}}$ represents the fusion function integrating both modalities. This fusion model integrates the tracking results of the RGB teacher model and the T student model, selecting the result with the higher confidence score in each frame to achieve robust tracking. For each $i$ $(i\in[1,I^{t}])$, $y_{t}[i]=f^{\text{T}}(x_{t}^{\text{T}})$, $\tilde{y}_{t}[j]=f^{\text{RGB}}(x_{t}^{\text{T}})$, and for the corresponding pair $j$ $(j\in[1,I^{t}])$ satisfying $y_{t}^{\text{ID}}[i]=\tilde{y}_{t}^{\text{ID}}[j]$, the final fusion result is determined as follows:
$\hat{y}_{t}[i]=\begin{cases}y_{t}[i]&\text{if }y_{t}^{\text{SCORE}}[i]>\tilde{y}_{t}^{\text{SCORE}}[j],\\ \tilde{y}_{t}[j]&\text{otherwise}.\end{cases}$ (6)
4 Evaluation Experiment
[…]
This dataset consists of a training sequence of 8,970 synchronized RGB thermal image pairs, and three image sequences obtained from both RGB and thermal cameras in the Bright scene, containing 1,081 frames, 1,049 frames, and 1,010 frames, respectively, and a single image sequence from both RGB and thermal cameras in the Dark scene, containing 666 frames. For performance comparison, the ground truth data is manually annotated with bounding boxes and person IDs.
Each tracking method is evaluated using Multiple Object Tracking Accuracy (MOTA) and ID F1 score (IDF1) to assess tracking performance, along with False Positive (FP), True Positive (TP), and False Negative (FN) to evaluate false detections [35]. ID Switches (IDSW) and Higher Order Tracking Accuracy (HOTA) measure trajectory reconstruction, with HOTA providing a global tracking assessment. Unlike MOTA, HOTA balances precision and recall by decomposing detection, assignment, and localization for a more detailed analysis (Table 1). These metrics gauge how effectively the proposed method tackles key challenges in Multiple Person Tracking (MPT): ID switching, detection failures, occlusion, and handling the appearance/disappearance of individuals. ID switching misassigns identities, detection failures disrupt continuity, and occlusion hides targets. Additionally, newly appearing or disappearing persons may be misidentified, reducing accuracy.
Table 1: Tracking Accuracy
Tracker MOTA IDF1 FP TP FN
RGB 38.3 56.3 26 1191 1822
T 93.5 97.0 113 2942 71
FUSION 92.4 96.6 136 2942 71
RGB: $f_{\theta^{\text{RGB}}}(x_{\text{Dark}}^{\text{T}})$, T: $f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$, FUSION: $f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}},x_{\text{Dark}}^{\text{T}})$.
Table 2: Trajectory Prediction Performance
Tracker HOTA IDSW
RGB 39.2 12
T 94.1 11
FUSION 93.4 21
$f_{\theta^{\text{RGB}}}(x_{\text{Full}}^{\text{T}})$ exhibited comparable performance in both Bright and Dark scenes. However, its performance did not reach that of $f_{\theta^{\text{RGB}}}(x_{\text{Bright}}^{\text{RGB}})$ in the Bright scene.
$f_{\theta^{\text{T}}}(x_{\text{Bright}}^{\text{RGB}})$+$f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$ achieved comparable performance in both Bright and Dark scenes.
$f_{\theta^{\text{T}}}(x_{\text{Full}}^{\text{T}})$ also demonstrated high performance comparable to $f_{\theta^{\text{T}}}(x_{\text{Bright}}^{\text{RGB}})$+$f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$ and, surprisingly, matched the performance of $f_{\theta^{\text{RGB}}}(x_{\text{Bright}}^{\text{RGB}})$ in the Bright scene (Fig. 3). The concern with this method is that it has the drawback of discarding high-resolution information inherent in RGB images.
$f_{\theta^{\text{RGB}}}(x_{\text{Bright}}^{\text{RGB}})$+$f_{\theta^{\text{RGB}}\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$ achieved slightly inferior results compared to $f_{\theta^{\text{RGB}}}(x_{\text{Bright}}^{\text{RGB}})$+$f_{\theta^{\text{T}}}(x_{\text{Dark}}^{\text{T}})$.
The above suggests that the most effective approach among all the combinations considered here is either switching between two trackers using a simple binary brightness classifier that requires RGB input or reusing the RGB tracker as a pseudo-annotation collector.
Figure 4 visualizes the results of the RGB tracker $f_{\theta^{\text{RGB}}}$ and the T tracker $f_{\theta^{\text{T}}}$ on T camera data. The RGB tracker $f_{\theta^{\text{RGB}}}$ failed to detect many targets, leading to tracking interruptions and false detections. In contrast, the T tracker $f_{\theta^{\text{T}}}$, benefiting from knowledge transfer, significantly improved tracking accuracy, reducing missed detections and enhancing ID stability. This is evident from the visually clearer reconstruction of multiple individuals’ boundaries and the effects of occlusion.
Figure 4: Visualization of $f_{\theta^{\text{T}}}(x^{\text{T}}_{\text{Bright}})$ (top) and $f_{\theta^{\text{RGB}}}(x^{\text{T}}_{\text{Bright}})$ (bottom). The three-dimensional $x$-$y$-$t$ trajectories are projected onto the xy plane and xt plane, shown in panels (a) and (b), respectively. In each panel, the three different columns represent different input image sequences, and different colors indicate different person IDs.
5 Conclusion
In this study, we tackled the cooperative MPT problem using both RGB and thermal (T) cameras placed in the same pose for the first time. As expected, the RGB tracker completely failed in the Dark scene’s RGB vision and also showed reduced performance in thermal vision. As a solution, we confirmed that learning the T tracker using pseudo-annotations from the RGB tracker is highly effective, achieving the best performance in both Bright and Dark scenes. We also investigated a fusion tracker that integrates both the RGB and T trackers, but surprisingly, its performance ranked second. This suggests that a cooperative approach dynamically switching between the two trackers is more effective than integration. In the future, we aim to apply the developed RGB-T cooperative MPT tracker to real-world robotic vision applications, including recognition, understanding, and reconstruction of individuals, occluders, and traversable areas.
Acknowledgements
Our work has been supported in part by JSPS KAKENHI Grant-in-Aid for Scientific Research (C) 20K12008 and 23K11270.